# The Pólya-Tchebotarev problem with semiclassical external fields

### v1

@@ 2.2 The class of admissible contours
Fix a potential $\varphi=\operatorname{Re}\Phi$ as in (2.5) with corresponding set of admissible angles $\Theta$ as described in (2.11) et seq., and a finite collection of points $\mathcal{C}$ as in (2.9). Let $\mathcal{P}(\mathcal{C})$ be a partition of $\mathcal{C}$, $\Psi:\mathcal{P}(\mathcal{C})\to 2^{\Theta}$ a function satisfying (2.15), (2.16) and (2.17), and $\mathcal{P}(\Theta)$ a partition of $\Theta$ subordinated to $\Psi$. We call $(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta))$ an admissible triple, and associate to it a collection of contours $\mathcal{T}(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta))$, referred to as admissible family as follows. Each set $\Gamma\in\mathcal{T}(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta))$ satisfies the conditions:
@@
    $\Gamma$ is a finite union of piecewise smooth Jordan arcs, possibly unbounded but satisfying (2.3), and with $\mathcal{C}\in\Gamma$.
@@
As a slight abuse of notation, we often write $\mathcal{T}$ and omit the triple $\mathcal{T}(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta))$ from our notation, and refer to $\mathcal{T}$ as a class of admissible contours rather than sets, to remind the reader that sets in $\mathcal{T}$ are obtained as unions of contours. Some expressions will be used to remove the technicality of the language of partitions given above:
@@
The weighted Pólya-Chebotarev’s problem for a finite set $\mathcal{C}$ can be obtained by setting $\mathcal{P}(\mathcal{C})=\{\mathcal{C}\}$, $\Psi(\mathcal{C})=\emptyset$ and $\mathcal{P}_{0}(\Theta_{\infty})=\emptyset$.
@@ 2.3 The main theorem
With the datum given by the external field $\varphi=\operatorname{Re}\Phi$ as in (2.5), a collection of fixed points $\mathcal{C}$ as in (2.9), and the class of contours $\mathcal{T}$ for an admissible triple $(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta))$, the associated Pólya-Chebotarev problem, or max-min energy problem asks for finding a set $\Gamma_{0}\in\mathcal{T}$ that maximizes the energy functional $\operatorname{I}^{\varphi}$ on $\mathcal{T}$, that is, finding $\Gamma_{0}$ such that
@@
    extends to a rational function on $\mathbb{C}$. Its simple poles belong to the set of fixed points $\mathcal{C}$, its double poles coincide with the logarithmic singularities of $\Phi$, and all its other poles have even order, always at least $4$, and coincide with the poles of the rational part of $\Phi$.
@@
As a consequence of the discussion in Examples 2.3 and 2.4, Theorem 2.5 includes both the Classical Pólya-Chebotarev Problem and the max-min considered in [14] as particular cases.
@@
$\frac{\partial}{\partial\eta_{+}}\left(\operatorname{U}^{\mu_{0}}(z)+\frac{1}{2}\varphi(z)\right)=\frac{\partial}{\partial\eta_{-}}\left(\operatorname{U}^{\mu_{0}}(z)+\frac{1}{2}\varphi(z)\right),$
on subarcs of $\operatorname{supp}\mu_{0}$, where $\eta_{\pm}$ are the two vectors normal to $\operatorname{supp}\mu_{0}$ at the point $z\in\operatorname{supp}\mu_{0}$. This property is known as the S-property in orthogonal polynomial theory.
@@ 2.4 Strategy
In [25] Rakhmanov presented an “ideal situation” under which a max-min weighted log energy problem admits a solution. However, such ideal situation is not always in place in many important concrete situations. In fact, an example of the type of problems for which his approach does not apply directly was mentioned already in [25], in the context of non-hermitian orthogonality with polynomial external fields. His approach was then adapted to this situation by Kuijlaars and the second-named author in [14], and our approach here follows their footsteps. When compared with [14], many of the technical lemmas need adaptation due to the fact that we now have to deal with singularities of $\Phi$ at finite points, not only in $\infty$. We highlight the distinctions along the way of our proofs. To explain Rakhmanov’s approach, define the hyperbolic metric on $\overline{\mathbb{C}}$ as the pull-back of the Euclidean metric on $T(\overline{\mathbb{C}})\subset\mathbb{R}^{3}$, where $T$ is the homeomorphism
@@
The Hausdorff metric above naturally induces a metric on the closed sets of $\mathbb{C}$, given by the distance between their closures on $\overline{\mathbb{C}}$. Now let $\mathcal{F}$ be a family of closed sets on $\mathbb{C}$. Assume in addition that the number of connected components of the elements of $\mathcal{F}$ is uniformly bounded by a finite value. Equip $\mathcal{F}$ with the Hausdorff metric just mentioned.
Theorem 2.6 ([25, Theorem 3.2]).
@@
If the class $\mathcal{F}$ is closed and $\operatorname{I}^{\varphi}$ is bounded from above by a finite value, then Theorem 2.6 implies the existence of a set $F_{0}$ maximizing $\operatorname{I}^{\varphi}$ on $\mathcal{F}$. This is not true for an admissible class of contours $\mathcal{T}$ as in Definition 2.2. Indeed, first of all, contours are not preserved under Hausdorff metric: one could construct a sequence space-filling curves converging to a set of strictly positive measure on $\mathbb{C}$, hence not a contour. But, in a more drastic issue, one could have a sequence of contours where each one extends to $\infty$ along given fixed sectors, but with a limiting contour that does not extend on the given sector, but in a different (neighboring) one. To overcome these issues, we need instead to work with a subclass $\mathcal{T}_{0}$ of $\mathcal{T}$, suitably chosen so that sets in the closure $\mathcal{F}_{0}=\overline{\mathcal{T}}_{0}$ of $\mathcal{T}_{0}$ in the Hausdorff metric have a better geometry than sets in the closure $\mathcal{F}=\overline{\mathcal{T}}$ of $\mathcal{T}$, in order to avoid the second issue mentioned above. Nevertheless, $\mathcal{F}_{0}$ still consists of sets more arbitrary than unions of contours. Thus, when we apply Theorem 2.6 to $\mathcal{F}_{0}$, we obtain a max-min set $F_{0}$ for the energy $\operatorname{I}^{\varphi}$ which is not necessarily a contour.
@@ 2.5 Structure for the rest of the paper
At last, we turn focus to the Pólya-Chebotarev Problem (2.18), and in Section 4 we combine the results of previous sections to prove Theorem 2.5.
@@ 3.1 Critical Measures
While studying the asymptotic behavior of zeroes of Stieltjes Polynomials, Martínez-Finkelshtein and Rakhmanov defined the notion of critical measures, measures that are critical points of a log energy functional. Since then this notion has found its way to other problems, for instance, in zero distribution of orthogonal polynomials and related families [19, 4], random matrix theory and other related models [20]. Recall that we are always assuming $\varphi=\operatorname{Re}\Phi$ of the form (2.5), with set of singularities $\mathcal{Z}=\{z_{1},\ldots,z_{n_{Q}},w_{1},\ldots,w_{n_{L}}\}$, and $\mathcal{C}=\{c_{1},\ldots,c_{n_{\mathcal{C}}}\}$ denotes a set of fixed points in $\mathbb{C}$. For this section, it is also convenient to write
$\Phi^{\prime}(z)=\frac{B(z)}{A(z)},\quad\text{with}\quad A(z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\prod_{j=1}^{n_{Q}}(z-z_{j})^{m_{j}+1}\prod_{j=1}^{n_{L}}(z-w_{j}),$ (3.1)
where $B$ is a polynomial of degree $\deg P+n_{Q}+n_{L}-1$ with $B(z)\neq 0$ for $z\in\mathcal{Z}$ whose leading coefficient is $N$ times the leading coefficient of $P$.
@@
= . . $\mathcal{H}=\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}\{h\in C_{c}^{2}(\mathbb{C});\;h(c)=0,c\in\mathcal{C},\text{ and }h\equiv 0\text{ in a neighborhood of }\mathcal{Z}\}.$ (3.2)
@@
The properties of $\mathcal{H}$ are tied to our later goal of perturbing a max-min contour, and using its extremal property within $\mathcal{T}$ to derive properties of its equilibrium measure, which will turn out to be critical. We impose $h$ to vanish at $\mathcal{C}$ so as that perturbations of the extremal contour indeed remain in $\mathcal{T}$. Later on, we will in fact show that critical measures cannot be supported near singularities of $\Phi$, but until we are able to prove this fact, we need to impose that members of the class $\mathcal{H}$ vanish near the singularities of $\Phi$, to avoid a blow-up of the energy when doing perturbations. Recall that $\mathcal{M}_{1}^{\varphi}(\mathbb{C})$ was previously defined in Section 2.1. It turns out that, within the class of test functions $\mathcal{H}$, directional derivatives of $\mu\in\mathcal{M}_{1}^{\varphi}(\mathbb{C})$ always exist and are somewhat explicit, as claimed by the next proposition. This proposition has appeared before, for instance, in [17, 14], but for completeness the proof here has to be adapted in virtue of the singularities of $\Phi$.
@@
where $O$ is a finite union of open discs $D_{\eta/2}(z_{j})$ in which $h$ is identically zero. If $x$ and $x+th(x)$ are not singularities of $\Phi$, we write
@@
where the term $r$ should be interpreted as an error. In particular, from this identity we obtain that $r(th(x);x)=0$ whenever $x\notin\operatorname{supp}h$. For sufficiently small $t$ one has that $x\in\mathbb{C}\setminus O$ implies $x+th(x)\in\mathbb{C}\setminus O$. Thus
@@
Now that we have a notion of derivative of $\operatorname{I}^{\varphi}$, the notion of critical point is natural.
@@
Martínez-Finkelshtein and Rakhmanov [17] showed that the Cauchy transform of a critical measure satisfies an algebraic equation, which in turn guarantees that the support of the measure lives in trajectories of a quadratic differential. Analogous results were obtained in [14, 19]. The main result of this section establishes the same connection between critical measures and algebraic equations. Following [17, 14], the main idea is to apply (3.7) to the so-called Shiffer variations $x\mapsto h(x)=H(x)/(x-\zeta)$, for each $\zeta\in\mathbb{C}$ fixed and $H(x)$ an appropriately chosen polynomial. However, to do so we need to extend the validity of Corollary 3.4 beyond the class of test functions $\mathcal{H}$, which we do so extending it in several steps. But before proceeding, we first establish an auxiliary approximation lemma.
@@
$\alpha_{n}(t)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\begin{cases}0,\quad t\leq\frac{1}{n},\\ (nx-1)^{3}(6n^{2}x^{2}-27nx+31),\quad\frac{1}{n}\leq t\leq\frac{2}{n}\\ 1,\quad\frac{2}{n}\leq t\end{cases}$
@@
Fix $M\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\deg B+\deg C$ points $x_{1},\ldots,x_{M}$ which are not singularities of $\Phi$, and a function $\chi\in C^{2}(\mathbb{C})$ which vanishes in a neighborhood of each point $x_{1},\ldots,x_{M}$, and is identically equal to $1$ in a for sufficiently large $z$. If $\mu\in\mathcal{M}_{1}^{\varphi}(\mathbb{C})$ is a critical measure, then Equation (3.7) remains valid for the choice
@@
$\psi_{n}(z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\beta_{n}(|z|)\prod_{w\in\mathcal{Z}}\alpha_{n}(|z-w|),\quad h_{n}(x)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\psi_{n}(z)h(z),\quad z\in\mathbb{C}.$ (3.9)
@@
Thus $\alpha_{n}\in C^{2}$. Analogously one sees that $\beta_{n}\in C^{2}_{c}$. This implies that $h_{n}\in\mathcal{H}$, so (3.7) is valid for it. From (3.8) we see that $\Phi^{\prime}h$ is continuous and bounded, and thus belongs to $L^{1}(\mathrm{d}\mu)$. Since $|\alpha_{n}|,|\beta_{n}|\leq 1$, we have $|\Phi^{\prime}h_{n}|\leq|\Phi^{\prime}h|$, and the Dominated Convergence Theorem implies that
@@
From a direct calculation,
@@
We view $h=h_{1}+\mathrm{i}h_{2}\simeq(h_{1},h_{2})$ as a function $\mathbb{R}^{2}\to\mathbb{R}^{2}$, with $h_{1}=\operatorname{Re}h,h_{2}=\operatorname{Im}h$ and Jacobian $Dh$. For $z\in\mathbb{C}$ and $(u,v)\in\mathbb{R}^{2}$, we differentiate from (3.9) and obtain
@@
Since $h_{1}(z)=\operatorname{Re}h(z)$, $h_{2}(z)=\operatorname{Im}h(z)$ and $h$ has a zero of order at least $1$ at each $w\in\mathcal{Z}$, we use (3.11) and conclude that the second term in the r.h.s. above is bounded. For the first term, note that $\|Dh(z)\|$ is bounded on compact sets since $h\in C^{2}$. Using that $h$ is holomorphic for sufficiently large $z$, we see that
@@
Therefore $\|Dh(z)\|=|h^{\prime}(z)|$ is bounded for sufficiently large $z$ and, by continuity, on the whole complex plane. In summary, we conclude from the inequality
$\left|\frac{h_{n}(x)-h_{n}(y)}{x-y}\right|\leq\sup_{z\in\mathbb{C}}\|Dh_{n}(z)\|$
that the quotient on the left-hand side is uniformly bounded and therefore, by the Dominated Convergence Theorem,
This concludes the proof. ∎
@@
When $\mathcal{Z}=\emptyset$, that is, when $\Phi$ is a polynomial, the result reduces to [14, Lemma 3.5]. The proof presented here is essentially the same, we briefly outline it and refer to [14] for more details.
@@
where $\epsilon_{n}$ is a sequence of positive numbers converging to zero. Note that $h_{n}\in H$ and thus (3.7) holds for it. Now, Equation (3.12) implies that $h\Phi^{\prime}$ is integrable with respect to $\mu$ and the dominated convergence theorem implies
@@
and since $g\in H$, there exists $M>0$ such that
@@
Now, Equation (3.12) implies that the right-hand side above is $\mu\times\mu$ integrable, and thus the Dominated convergence theorem implies
@@
    Each pole $z_{j}$ of order $m_{j}$ of $\Phi$ is a pole of order $22m_{j}+$ 2 of $R$.
@@
    Each logarithmic singularity $w_{i}$ of $\Phi$ is a double pole of $R$.
@@
The function $h$ satisfies the conditions of Corollary 3.8, and therefore Equation (3.7) is valid for it. Our next goal is to rewrite both sides of (3.7) for this $h$. We start by writing
@@
where $P_{3}$ is a polynomial of degree at most $M-2$ in each of its variables $x,y,z$. This shows that $D_{2}$ is a rational function as well, with only simple poles, all contained in the set $\{x_{1},\ldots,x_{M-1}\}$, and furthermore
@@
The key point here is that a critical measure $\mu$ satisfies an algebraic equation of the form (3.14). Furthermore, by definition of the class $\mathcal{M}_{1}^{\varphi}(\mathbb{C})$, the critical measure $\mu$ has finite log energy and therefore it cannot have mass points. Once these properties are known, the proof follows in the same way as in the proof of the analogous result [14, Proposition 3.8], so we skip the details. ∎
@@
Let $\mu$ be a critical measure. The support of $\mu$ is a compact set made of a finite union of analytic arcs, that do not contain any of the poles of $\Phi$.
@@
  1. (1)
    this union is over finitely many arcs,

### v2

@@ 2.2 The class of admissible contours
Fix a potential $\varphi=\operatorname{Re}\Phi$ as in (2.5) with corresponding set of admissible angles $\Theta$ as described in (2.11) et seq., and a finite collection of points $\mathcal{C}$ as in (2.9).
Let $\mathcal{P}(\mathcal{C})$ be a partition of $\mathcal{C}$, $\Psi:\mathcal{P}(\mathcal{C})\to 2^{\Theta}$ a function satisfying (2.15), (2.16) and (2.17), and $\mathcal{P}(\Theta)$ a partition of $\Theta$ subordinated to $\Psi$. Assume that $\mathcal{C}\neq\emptyset$ or $\mathcal{P}_{0}(\Theta)\neq\emptyset$. We call $(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta))$ an admissible triple, and associate to it a collection of contours $\mathcal{T}(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta))$, referred to as admissible family as follows. Each set $\Gamma\in\mathcal{T}(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta))$ satisfies the conditions:
@@
    $\Gamma$ is a finite union of piecewise smooth Jordan arcs, possibly unbounded but satisfying (2.3), and with $\mathcal{C}\subset\Gamma$.
@@
As a slight abuse of notation, we often write $\mathcal{T}$ and omit the triple $\mathcal{T}(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta))$ from our notation, and refer to $\mathcal{T}$ as a class of admissible contours rather than sets, to remind the reader that sets in $\mathcal{T}$ are obtained as unions of contours.
Some expressions will be used to remove the technicality of the language of partitions given above:
@@
The weighted Pólya-Tchebotarev’s problem for a finite set $\mathcal{C}$ can be obtained by setting $\mathcal{P}(\mathcal{C})=\{\mathcal{C}\}$, $\Psi(\mathcal{C})=\emptyset$ and $\mathcal{P}_{0}(\Theta_{\infty})=\emptyset$.
@@ 2.3 The main theorem
With the data given by the external field $\varphi=\operatorname{Re}\Phi$ as in (2.5), a collection of fixed points $\mathcal{C}$ as in (2.9), and the class of contours $\mathcal{T}$ for an admissible triple $(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta))$, the associated Pólya-Tchebotarev problem, or max-min energy problem asks for finding a set $\Gamma_{0}\in\mathcal{T}$ that maximizes the energy functional $\operatorname{I}^{\varphi}$ on $\mathcal{T}$, that is, finding $\Gamma_{0}$ such that
@@
    extends to a rational function on $\mathbb{C}$. Its simple poles belong to the set of fixed points $\mathcal{C}$, its double poles coincide with points $w_{1},\ldots,w_{n_{L}}$, and all its other poles have even order, always at least $4$, and coincide with the poles of the rational part of $\Phi$.
@@
As a consequence of the discussion in Examples 2.3 and 2.4, Theorem 2.5 includes both the Classical Pólya-Tchebotarev Problem and the max-min considered in [18] as particular cases.
@@
$\frac{\partial}{\partial\eta_{+}}\left(\operatorname{U}^{\mu_{0}}(z)+\varphi(z)\right)=\frac{\partial}{\partial\eta_{-}}\left(\operatorname{U}^{\mu_{0}}(z)+\varphi(z)\right),$
on subarcs of $\operatorname{supp}\mu_{0}$, where $\eta_{\pm}$ are the two unit vectors normal to $\operatorname{supp}\mu_{0}$ at the point $z\in\operatorname{supp}\mu_{0}$. This property is known as the S-property in orthogonal polynomial theory.
@@ 2.4 Strategy
In [30] Rakhmanov presented an “ideal situation” under which a max-min weighted log energy problem admits a solution. However, such ideal situation is not always in place in many important concrete situations. In fact, an example of the type of problems for which his approach does not apply directly was mentioned already in [30], in the context of non-hermitian orthogonality with polynomial external fields. His approach was then adapted to this situation by Kuijlaars and the second-named author in [18], and our approach here follows their footsteps. When compared with [18], many of the technical lemmas need adaptation due to the fact that we now have to deal with singularities of $\Phi$ at finite points, not only in $\infty$. We highlight the distinctions along the way of our proofs.
To explain Rakhmanov’s approach, define the hyperbolic metric on $\overline{\mathbb{C}}$ as the pull-back of the Euclidean metric on $T(\overline{\mathbb{C}})\subset\mathbb{R}^{3}$, where $T$ is the homeomorphism
@@
The Hausdorff metric above naturally induces a metric on the closed sets of $\mathbb{C}$, given by the distance between their closures on $\overline{\mathbb{C}}$.
Now let $\mathcal{F}$ be a family of closed sets on $\mathbb{C}$. Assume in addition that the number of connected components of the elements of $\mathcal{F}$ is uniformly bounded by a finite value. Equip $\mathcal{F}$ with the Hausdorff metric just mentioned.
Theorem 2.6 ([30, Theorem 3.2]).
@@
If the class $\mathcal{F}$ is closed and $\operatorname{I}^{\varphi}$ is bounded from above by a finite value, then Theorem 2.6 implies the existence of a set $F_{0}$ maximizing $\operatorname{I}^{\varphi}$ on $\mathcal{F}$. This is not true for an admissible class of contours $\mathcal{T}$ as in Definition 2.2. Indeed, first of all, contours are not preserved under Hausdorff metric: one could construct a sequence of space-filling curves converging to a set of strictly positive measure on $\mathbb{C}$, hence not a contour. But, in a more drastic issue, one could have a sequence of contours where each one extends to $\infty$ along given fixed sectors, but with a limiting contour that does not extend on the given sector, but in a different (neighboring) one.
To overcome these issues, we need instead to work with a subclass $\mathcal{T}_{0}$ of $\mathcal{T}$, suitably chosen so that sets in the closure $\mathcal{F}_{0}=\overline{\mathcal{T}}_{0}$ of $\mathcal{T}_{0}$ in the Hausdorff metric have a better geometry than sets in the closure $\mathcal{F}=\overline{\mathcal{T}}$ of $\mathcal{T}$, in order to avoid the second issue mentioned above.
Nevertheless, $\mathcal{F}_{0}$ still consists of sets more arbitrary than unions of contours. Thus, when we apply Theorem 2.6 to $\mathcal{F}_{0}$, we obtain a max-min set $F_{0}$ for the energy $\operatorname{I}^{\varphi}$ which is not necessarily a contour.
@@ 2.5 Structure for the rest of the paper
At last, we turn focus to the Pólya-Tchebotarev Problem (2.18), and in Section 4 we combine the results of previous sections to prove Theorem 2.5.
@@ 3.1 Critical Measures
While studying the asymptotic behavior of zeroes of Stieltjes Polynomials, Martínez-Finkelshtein and Rakhmanov defined the notion of critical measures, measures that are critical points of a log energy functional. Since then this notion has found its way to other problems, for instance, in zero distribution of orthogonal polynomials and related families [23, 5], random matrix theory and other related models [24].
We always assume $\varphi=\operatorname{Re}\Phi$, where $\Phi$ was defined in Equation (2.5). We also fix the notation $\mathcal{Z}=\{z_{1},\ldots,z_{n_{Q}},w_{1},\ldots,w_{n_{L}}\}$ for the set of singularities, $\mathcal{W}=\{w_{1},\ldots,w_{n_{L}}\}\setminus\mathcal{Z}$ and $\mathcal{C}=\{c_{1},\ldots,c_{n_{\mathcal{C}}}\}$ for the set of fixed points. We also consider the polynomial
$C(z)=\prod_{i=1}^{n_{C}}(z-c_{i})$
with $C\equiv 1$ when $\mathcal{C}=\emptyset$.
For this section, it is also convenient to write
$\Phi^{\prime}(z)=\frac{B(z)}{A(z)},\quad\text{with}\quad A(z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\prod_{j=1}^{n_{Q}}(z-z_{j})^{m_{j}+1}\prod_{w\in\mathcal{W}}(z-w_{j}),$ (3.1)
where $B$ is a polynomial of degree $\deg P+n_{Q}+|\mathcal{W}|-1$ with $B(z)\neq 0$ for $z\in\mathcal{W}$.
@@
. . =$\mathcal{H}\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\{h\in C_{c}^{2}(\mathbb{C});\;h(c)=0,c\in\mathcal{C},\text{ and }h\equiv 0\text{ in a neighborhood of }\mathcal{Z}\}.$ (3.2)
@@
The properties of $\mathcal{H}$ are tied to our later goal of perturbing a max-min contour, and using its extremal property within $\mathcal{T}$ to derive properties of its equilibrium measure, which will turn out to be critical. We impose $h$ to vanish at $\mathcal{C}$ so as that perturbations of the extremal contour indeed remain in $\mathcal{T}$. Later on, we will in fact show that critical measures cannot be supported near singularities of $\Phi$, but until we are able to prove this fact, we need to impose that members of the class $\mathcal{H}$ vanish near the singularities of $\Phi$, to avoid a blow-up of the energy when doing perturbations.
Recall that $\mathcal{M}_{1}^{\varphi}(\mathbb{C})$ was previously defined in Section 2.1. It turns out that, within the class of test functions $\mathcal{H}$, directional derivatives of $\mu\in\mathcal{M}_{1}^{\varphi}(\mathbb{C})$ always exist and are somewhat explicit, as claimed by the next proposition. This proposition has appeared before, for instance, in [21, 18], but for completeness the proof here has to be adapted in virtue of the singularities of $\Phi$.
@@
where $O$ is a finite union of open discs $D_{\eta/2}(z_{j})$ in which $h$ is identically zero.
If $x$ and $x+th(x)$ are not singularities of $\Phi$, we write
@@
where the term $r$ should be interpreted as an error. In particular, from this identity we obtain that $r(th(x);x)=0$ whenever $x\notin\operatorname{supp}h$.
For sufficiently small $t$ one has that $x\in\mathbb{C}\setminus O$ implies $x+th(x)\in\mathbb{C}\setminus O$. Thus
@@
Now that we have a notion of derivative of $\operatorname{I}^{\varphi}$, the notion of critical measure is natural.
@@
Martínez-Finkelshtein and Rakhmanov [21] showed that the Cauchy transform of a critical measure satisfies an algebraic equation, which in turn guarantees that the support of the measure lives in trajectories of a quadratic differential. Analogous results were obtained in [18, 23]. The main result of this section establishes the same connection between critical measures and algebraic equations. Following [21, 18], the main idea is to apply (3.7) to the so-called Schiffer variations $x\mapsto h(x)=H(x)/(x-\zeta)$, for each $\zeta\in\mathbb{C}$ fixed and $H(x)$ an appropriately chosen polynomial. However, to do so we need to extend the validity of Corollary 3.4 beyond the class of test functions $\mathcal{H}$, which we do so extending it in several steps. But before proceeding, we first establish an auxiliary approximation lemma.
@@
$\alpha_{n}(t)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\begin{cases}0,\quad t\leq\frac{1}{n},\\ (nt-1)^{3}(6n^{2}t^{2}-27nt+31),\quad\frac{1}{n}\leq t\leq\frac{2}{n}\\ 1,\quad\frac{2}{n}\leq t\end{cases}$
@@
Fix $M\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\deg B+\deg C$ points $x_{1},\ldots,x_{M}$ which are not singularities of $\Phi$, and a function $\chi\in C^{2}(\mathbb{C})$ which vanishes in a neighborhood of each point $x_{1},\ldots,x_{M}$, and is identically equal to $1$ in a neighborhood of $\infty$. If $\mu\in\mathcal{M}_{1}^{\varphi}(\mathbb{C})$ is a critical measure, then Equation (3.7) remains valid for the choice
@@
$\psi_{n}(z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\beta_{n}(|z|)\prod_{w\in\mathcal{Z}}\alpha_{n}(|z-w|),\quad h_{n}(z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\psi_{n}(z)h(z),\quad z\in\mathbb{C}.$ (3.9)
@@
Thus $\alpha_{n}\in C^{2}$. Analogously one sees that $\beta_{n}\in C^{2}_{c}$. This implies that $h_{n}\in\mathcal{H}$, so (3.7) is valid for it.
From (3.8) and $M=\deg B+\deg C$, we see that $\Phi^{\prime}h$ is continuous and bounded, and thus belongs to $L^{1}(\mathrm{d}\mu)$. Since $|\alpha_{n}|,|\beta_{n}|\leq 1$, we have $|\Phi^{\prime}h_{n}|\leq|\Phi^{\prime}h|$, and the Dominated Convergence Theorem implies that
@@
The Mean Value Inequality implies that
$\left|\frac{h_{n}(x)-h_{n}(y)}{x-y}\right|\leq\sup_{z\in\mathbb{C}}\|Dh_{n}(z)\|.$
Thus, provided that we show that the right-hand side is finite and bounded uniformly in $n\in\mathbb{N}$, the Dominated Convergence Theorem gives
We now verify that $\|Dh_{n}(z)\|$ is bounded uniformly in $n$. From a direct calculation,
@@
We view $h=h_{1}+\mathrm{i}h_{2}\simeq(h_{1},h_{2})$ as a function $\mathbb{R}^{2}\to\mathbb{R}^{2}$, with $h_{1}=\operatorname{Re}h,h_{2}=\operatorname{Im}h$ and Jacobian $Dh$. For $z\in\mathbb{C}$ and $(u,v)\in\mathbb{R}^{2}$, we differentiate (3.9) and obtain
@@
Since $h_{1}(z)=\operatorname{Re}h(z)$, $h_{2}(z)=\operatorname{Im}h(z)$ and $h$ has a zero of order at least $1$ at each $w\in\mathcal{Z}$, we use (3.11) and conclude that the second term in the r.h.s. above is bounded.
For the first term, note that $\|Dh(z)\|$ is bounded on compact sets since $h\in C^{2}$. Using that $h$ is holomorphic for sufficiently large $z$, we see that
@@
Therefore $\|Dh(z)\|=|h^{\prime}(z)|$ is bounded a neighborhood of infinity and, thus, on the whole complex plane. This concludes the proof. ∎
@@
When $\mathcal{Z}=\emptyset$, that is, when $\Phi$ is a polynomial, the result reduces to [18, Lemma 3.5]. The proof presented here is essentially the same, we briefly outline it and refer to [18] for more details.
@@
where $\epsilon_{n}$ is a sequence of positive numbers converging to zero. Note that $h_{n}\in\mathcal{H}$ and thus (3.7) holds for it. Now, Inequality (3.12) implies that $h\Phi^{\prime}$ is integrable with respect to $\mu$ and the dominated convergence theorem implies
@@
and since $g\in\mathcal{H}$, there exists $M>0$ such that
@@
Now, Inequality (3.12) implies that the right-hand side above is $\mu\times\mu$ integrable, and thus the Dominated convergence theorem implies
@@
    Each $z_{j}$ of order $m_{j}$ of $\Phi$ is a pole of order $22m_{j}+$ 2 of $R$.
@@
    Each $w_{i}\in\mathcal{W}$ is a double pole of $R$.
@@
The function $h$ satisfies the conditions of Corollary 3.8, and therefore Equation (3.7) is valid for it. Our next goal is to rewrite both sides of (3.7) for this $h$.
We start by writing
@@
where $P_{2}$ is a polynomial of degree at most $M-2$ in each of its variables $x,y,z$. This shows that $D_{2}$ is a rational function as well, with only simple poles, all contained in the set $\{x_{1},\ldots,x_{M-1}\}$, and furthermore
@@
The key point here is that a critical measure $\mu$ satisfies an algebraic equation of the form (3.14). Furthermore, by definition of the class $\mathcal{M}_{1}^{\varphi}(\mathbb{C})$, the critical measure $\mu$ has finite log energy and therefore it cannot have mass points. Once these properties are known, the proof follows in the same way as in the proof of the analogous result [18, Proposition 3.8], so we skip the details. ∎
@@
Let $\mu$ be a critical measure. The support of $\mu$ is a compact set made of a union of analytic arcs, that do not contain any of the zeroes of $Q$.